# Take Fake as Real: Realistic-like Robust Black-box Adversarial Attack to Evade AIGC Detection

### v1

@@ Abstract
The security of AI-generated content (AIGC) detection based on GANs and diffusion models is closely related to the credibility of multimedia content. Malicious adversarial attacks can evade these developing AIGC detection. However, most existing adversarial attacks focus only on GAN-generated facial images detection, struggle to be effective on multi-class natural images and diffusion-based detectors, and exhibit poor invisibility. To fill this gap, we first conduct an in-depth analysis of the vulnerability of AIGC detectors and discover the feature that detectors vary in vulnerability to different post-processing. Then, considering the uncertainty of detectors in real-world scenarios, and based on the discovery, we propose a Realistic-like Robust Black-box Adversarial attack (R2BA) with post-processing fusion optimization. Unlike typical perturbations, R2BA uses real-world post-processing, i.e., Gaussian blur, JPEG compression, Gaussian noise and light spot to generate adversarial examples. Specifically, we use a stochastic particle swarm algorithm with inertia decay to optimize post-processing fusion intensity and explore the detector’s decision boundary. Guided by the detector’s fake probability, R2BA enhances/weakens the detector-vulnerable/detector-robust post-processing intensity to strike a balance between adversariality and invisibility. Extensive experiments on popular/commercial AIGC detectors and datasets demonstrate that R2BA exhibits impressive anti-detection performance, excellent invisibility, and strong robustness in GAN-based and diffusion-based cases. Compared to state-of-the-art white-box and black-box attacks, R2BA shows significant improvements of 15% and 21% in anti-detection performance under the original and robust scenario respectively, offering valuable insights for the security of AIGC detection in real-world applications.
@@
Adversarial attack, AIGC detection, Post-processing, Particle swarm optimization
@@ I Introduction
With the rapid development of GAN- and diffusion-based generation technologies[1], AI-generated content (AIGC) detection has become an increasingly popular research topic[2],[3],[4]. The security of AIGC detection is closely linked to the credibility of multimedia content[5]. If the security of AIGC detection technology is not guaranteed, multimedia content will face a serious trust crisis. However, AIGC detection technologies are still in an immature stage: Many studies[6],[7],[8],[9] have shown that GAN-based image forgery detectors are vulnerable to malicious adversarial attacks, resulting in detection failures. On the other hand, the widespread use of diffusion models makes security research on diffusion-based detection increasingly urgent. However, the lack of robustness research on this detection method[2],[4],[10],[11],[12], combined with the fact that the generative mechanisms of diffusion models are completely different from those of GANs[13],[14], leaves the security of diffusion-based AIGC detection uncertain. Therefore, to enhance the security of AIGC, it is crucial to focus on the robustness of its detection against adversarial attacks, particularly those diffusion-based methods.
@@
Previous researches[9],[17],[18],[19],[20] demonstrate that adversarial examples can be generated by employing natural degradation as an attack method. These methods generate adversarial examples by simulating the degradation of image quality in the real world, which is more meaningful for studying the robustness of the AIGC detectors, as these post-processing techniques, like JPEG compression and Gaussian blur, are common on social networks. In other words, if a detector classifies such post-processed AIGC images as real, it suggests that the detector will face similar performance degradation issues in real-world applications. Therefore, employing post-processing techniques to generate adversarial samples is both feasible and of significant practical value.
In order to design a realistic-like adversarial attack applicable to natural images, as shown in Fig. 1, we select Gaussian blur, JPEG compression, Gaussian noise, and light spot as attacks to generate adversarial examples. These 4 techniques have been proven to attack the frequency, spatial, and statistical domains of detectors, respectively[9],[17],[19],[20]. And the pre-experiment in Fig. 2 reveal the discovery that detectors vary in vulnerability to different post-processing. Based on the discovery, and considering the uncertainty of detectors in real-world scenarios, we propose R2BA, a realistic-like robust black-box adversarial attack based on post-processing fusion optimization. During the optimization process, R2BA first randomly generates multiple post-processed adversarial examples to explore the detector’s complex decision boundary. Then, guided by the fake probability output of the AIGC detector, R2BA continuously updates the post-processing intensity of these adversarial examples, driving them toward and eventually across the decision boundary of the detector, thereby evading AIGC detection.
R2BA solves 3 problems existed in previous work:
Applicability to AIGC detectors. Based on the discovery that detectors vary in vulnerability to different post-processing, we integrate 4 widely-used post-processing techniques to launch a comprehensive attack, ensuring all detectors are successfully attacked by R2BA, resulting in significant performance degradation.
@@
  • We propose R2BA, a novel realistic-like robust black-box adversarial attack, which is the first method with significant practical value, capable of evading AIGC detectors on natural images. It demonstrates applicability to black-box detectors, with impressive anti-detection performance, excellent image quality, and strong robustness.
  • We propose a post-processing fusion optimization to achieve a trade-off between adversariality and invisibility. Through a stochastic particle swarm optimization algorithm with inertial decay, and oriented to the detector vulnerability to different post-processing, this optimization realizes the dynamic adjustment of the fusion intensity.
  • We conduct extensive experiments on 3 popular open-source AIGC image datasets, validating the effective anti-detection performance of R2BA in evading state-of-the-art GAN-based and diffusion-based AIGC detectors, as well as commercial AIGC detectors, in the original, robust and real-world scenarios.
The remainder of this paper is organized as follows. Section II reviews related work. Section III presents the specific implementation of the proposed method. The experiments and analyses are described in Section IV. Finally, we provide a brief conclusion in Section V.
@@ II-A AI-generated Image Forgery Detection
AI-generated image forgery detection is actually a Fake-Real binary classification problem. Here, Fake refers to AIGC, which means the image is generated by AI, specifically through techniques like GANs or diffusion models. While Real refers to UGC, meaning that the image is a natural image or user-generated. The challenge of AIGC detection lies in extracting subtle forgery signals from AI-generated images and making accurate distinctions. According to different generation technologies, AIGC detection techniques can be divided into GAN based and diffusion based methods.
GAN-based forgery detection. Image generation based on GAN architecture[21],[22],[23] consists of generative and discriminative models. The goal of this architecture is to train the generative model to exactly fit the real data distribution, so that the discriminative model is difficult to distinguish. To prevent this technique from being misused, several GAN-based forgery detection techniques[10],[11],[24],[25],[26] have been proposed. These works recognize AI-generated images by extracting specific features and use data-driven approaches to guide the model in learning the differences in features across the spatial, frequency, and statistical domains between AI-generated and natural images.
@@ II-B Adversarial Attacks for Anti-detection
Black-box attack. Black-box adversarial attacks do not require access to the inner workings of the model, and efficient attacks can be achieved by only obtaining the model output. There are three main types of black-box attacks: Algorithms based on gradient estimation[7] and agent models[30] are limited due to their reliance on fitting the target. When applied in practice, query-based methods might be the best solution. However, when facing the unpredictable decision boundaries caused by multi-class natural images[31], a single adversarial example requires significant effort to locate the decision boundary. Some methods based on random pixel updates[16],[32] are effective in anti-detection, but tend to cause significant perturbations, which allows these adversarial examples to be directly classified as forged images by the human eye, even without a detector. Another method involving attribute editing[15], although optimizing multiple adversarial examples simultaneously, still face the challenge of selecting the specific attribute across the diverse natural images and requires appropriate attribute editing tools.
Based on the above analysis, existing adversarial attacks for anti-detection have limitations in image classes (they only work on facial datasets) and detectors (primarily GAN-based detection). Additionally, they are often difficult to strike a balance between the anti-detection effectiveness of adversarial examples and image quality. Moreover, in practical applications, real-world platform image processing may erase the adversarial perturbations, leading to the failure of anti-detection. To address the issues in existing works, we conduct comprehensive experiments on the security of GAN-based and diffusion-based AIGC detectors using natural image datasets with multiple classes, and we thoroughly analyze and compare various anti-detection performance metrics of our method with existing methods in original, robust and real-world scenarios.
Figure 2: Average performance degradation. The values represent the attack success rate of different post-processing against AIGC detectors. Darker colors indicate greater vulnerability of detectors. Attacks used include Gaussian blur, JPEG compression, Gaussian noise, light spot, and sequential fusion attacks.
@@ III-A Motivation
Traditional adversarial examples[6],[7],[16],[32] generated by gradient-based attacks or random pixel modifications attacks, are both meaningless and ineffective because their poor invisibility makes them difficult to apply in real-world scenarios. They can be easily identified as fake by the human eye, without the need for a detector. In contrast, using post-processing techniques commonly employed in real-life scenarios to create adversarial examples makes them look realistic and match real natural image patterns. Furthermore, if a detector classifies such post-processed AIGC images as real, this suggests that the detector will face similar performance degradation in real-world applications.
@@
Inspired by previous works[9],[18],[20], we select Gaussian blur, JPEG compression, Gaussian noise and light spot as attacks, which have been proven to mislead detectors through frequency, spatial, and statistical domains respectively. Pre-experiment we conduct in Fig. 2 further confirm that these post-processing methods lead to varying degrees of performance degradation of detectors. We use Gaussian blur with a radius of $5$ and deviation of $1.5$, JPEG compression with an intensity of $80$, Gaussian noise with standard deviation of $7/255$, and light spot with an intensity of $1.8$ and random radius. The datasets[4],[24] and detectors[4],[10],[11],[24],[25] used in Fig. 2 are described in Section IV.
@@
The results in Fig. 2 demonstrate that detectors vary in vulnerability to different post-processing. For example, on the CNNSpot[24] dataset, the CNNSpot[24] detector is highly vulnerable to Gaussian noise but more robust to JPEG compression (GN=0.12, JG=0.00), while the FatFormer[11] detector exhibits the opposite behavior (GN=0.01, JG=0.23). Interestingly, the effectiveness of post-processing fusion attacks does not always exceed that of individual post-processing on different detectors. On the DRCT[4] dataset, the attack performance of Gaussian blur + JPEG compression on the FatFormer[11] is better than that of Gaussian blur alone (GB+JG=0.87, GB=0.18), whereas the opposite (GB+JG=0.02, GB=0.12) is true for the UnivFD[10]. Therefore, optimizing the intensity of post-processing fusion is challenging, but necessary.
@@
Therefore, considering the uncertainty of detectors in real-world scenarios, we adopt a black-box attack to generate adversarial examples. To attack different types of detectors, we fuse these four post-processing and define the attack order as step $1$-$4$ in Fig. 1. To solve the challenge of optimizing the fusion intensity, we use the detector’s fake probability as guidance and apply stochastic particle swarm optimization (PSO)[33] with inertia decay to optimize the fusion intensity. Compared to other optimization algorithms, such as evolutionary algorithms and simulated annealing, stochastic PSO with inertia decay outperforms them in its ability to effectively perform high-dimensional global searches[34]. This method can quickly locate the global decision boundary and simultaneously optimize fusion intensity to balance invisibility and adversariality. All things considered, we propose a post-processing fusion optimization-based, realistic-like robust black-box adversarial attack method, R2BA, to study the security of AIGC detectors in real-world applications.
III-B Visualisation
@@
Figure 3: The pipeline of R2BA. It contains 4 modules and a main optimization process (Initialization-Optimization iteration-Select). R2BA uses N randomly generated post-processing parameters as particle positions, optimizing their velocities and positions to balance the invisibility and adversariality. A successful attack occurs when a particle crosses the detector’s decision boundary, and R2BA outputs the adversarial example with the highest SSIM value among the successful ones. The notations’ definitions here are the same as in Table I.
@@
Component description. In Fig 3, we illustrate the pipeline of R2BA. It mainly consists of four modules: the post-processing module, fitness calculation, velocity/position update, and the swarm optimization module. The decision boundary here refers to the fake threshold ($0.5$) of the detector, which typically classifies an image with a fake probability greater than $0.5$ as fake and less than $0.5$ as real.
@@
1) Post-processing Module: The red circle represents an AIGC image and the blue represent its post-processed images (adversarial particles). The particle positions represent the post-processing fusion intensity. We can balance invisibility and adversariality by optimizing the particle positions.
@@
2) Fitness Calculation: The post-processed image is fed into the detector to obtain its fake probability, which serves as its particle’s fitness. A smaller fitness value indicates a more effective adversarial example. The individual best of each particle has the smallest fitness in the optimization process, i.e., the lowest fake probability achieved so far. The global best is the particle with the smallest fitness among all individual bests. The global best particle is most likely to cross the decision boundary, as it has the lowest fake probability.
@@
3) Velocity / Position Update: Velocity $V_{i}^{k+1}$ updates are influenced by the particle’s individual best $p_{best}$, global best $P_{best}$, and inertia $w^{k}$. While $p_{i}^{k+1}$ position updates depend on the current $p_{i}^{k}$ position and velocity $V_{i}^{k+1}$. Obviously, influenced by the individual best and global best, the particles will move towards positions with a lower fake probability, continuously approaching the decision boundary.
@@
Optimization Iteration. In the first iteration, to cover as much of the example domain space as possible, R2BA randomly assigns $N$ post-processing fusion intensity values as the initial positions of $N$ particles. Stochastic generation helps quickly locate the detector’s most vulnerable decision boundary, enabling efficient global search and addressing the challenges posed by diverse decision boundaries in multi-class natural images. In subsequent iterations, take the $k$-th iteration as an example, guided by the fitness, the particle velocity gradually points towards the decision boundary, with the swarm becoming more concentrated. In the final (or successful) iteration, some particles cross the boundary, causing the detector to misclassify them as real images. R2BA then output the adversarial example with the highest SSIM among all the particles with successful attacks.
@@
The above description outlines a typical iterative process. However, in practice, AIGC images may be located near the decision boundary, causing the attack to succeed in the first iteration. If the images are far from the decision boundary, the attack may fail to cross it within the range defined by the post-processing intensity settings. Experiments show that the former case is more common, while the latter is relatively rare.

### v2

@@ Abstract
The security of AI-generated content (AIGC) detection is crucial for ensuring multimedia content credibility. To enhance detector security, research on adversarial attacks has become essential. However, most existing adversarial attacks focus only on GAN-generated facial images detection, struggle to be effective on multi-class natural images and diffusion-based detectors, and exhibit poor invisibility. To fill this gap, we first conduct an in-depth analysis of the vulnerability of AIGC detectors and discover the feature that detectors vary in vulnerability to different post-processing. Then, considering that the detector is agnostic in real-world scenarios and given this discovery, we propose a Realistic-like Robust Black-box Adversarial attack (R2BA) with post-processing fusion optimization. Unlike typical perturbations, R2BA uses real-world post-processing, i.e., Gaussian blur, JPEG compression, Gaussian noise and light spot to generate adversarial examples. Specifically, we use a stochastic particle swarm algorithm with inertia decay to optimize post-processing fusion intensity and explore the detector’s decision boundary. Guided by the detector’s fake probability, R2BA enhances/weakens the detector-vulnerable/detector-robust post-processing intensity to strike a balance between adversariality and invisibility. Extensive experiments on popular/commercial AIGC detectors and datasets demonstrate that R2BA exhibits impressive anti-detection performance, excellent invisibility, and strong robustness in GAN-based and diffusion-based cases. Compared to state-of-the-art white-box and black-box attacks, R2BA shows significant improvements of 15%–72% and 21%–47% in anti-detection performance under the original and robust scenario respectively, offering valuable insights for the security of AIGC detection in real-world applications.
@@
Adversarial attack, AIGC detection, Image processing, Particle swarm optimization
@@ I Introduction
With the rapid development of GAN- and diffusion-based generation technologies[1], AI-generated content (AIGC) detection has become a popular research topic[2],[3],[4]. The security of state-of-the-art (SOTA) AIGC detection is crucial for ensuring multimedia content credibility[5]. Without secure AIGC detection technologies, multimedia content could face a serious trust crisis. However, AIGC detection is still in an immature stages: Many studies[6],[7],[8],[9] have shown that GAN-based detectors are vulnerable to adversarial attacks, leading to detection failures. On the other hand, the widespread use of diffusion models makes security research on this detection method more urgent. However, the lack of robustness research[2],[4],[10],[11],[12], combined with the distinct generative mechanisms of diffusion models compared to GANs[13],[14], leaves the security of diffusion-based AIGC detection uncertain. Therefore, to improve the security of AIGC detection, it is crucial to focus on the research of these technologies’ robustness against adversarial attacks, particularly for diffusion-based detectors.
@@
Previous researches[9],[17],[18],[19],[20] demonstrate that adversarial examples can be generated by employing natural images degradation as an attack method. These methods generate adversarial examples by simulating the degradation of image quality in the real world, which is more meaningful for studying the robustness of AIGC detectors, as post-processing techniques, like JPEG compression and Gaussian blur, are common on social networks. In other words, if a detector classifies such post-processed AIGC images as real, it suggests that the detector will face similar performance degradation issues in real-world applications. Therefore, employing post-processing techniques to generate adversarial samples is both feasible and of significant practical value.
In order to design a realistic-like adversarial attack applicable to natural images, as shown in Fig. 1, we select Gaussian blur, JPEG compression, Gaussian noise, and light spot as attacks to generate adversarial examples. These 4 techniques have been proven to attack the frequency, spatial, and statistical domains of detectors, respectively[9],[17],[19],[20]. And the pre-experiment we conduct in Fig. 2 reveals the discovery that detectors vary in vulnerability to different post-processing. Based on the discovery, and considering that the detector is agnostic in real-world scenarios, we propose R2BA, a realistic-like robust black-box adversarial attack based on post-processing fusion optimization. During the optimization process, R2BA first randomly generates multiple post-processed adversarial examples to explore the detector’s complex decision boundary. Then, guided by the fake probability output of the AIGC detector, R2BA continuously updates the post-processing intensity of these adversarial examples, driving them toward and eventually across the decision boundary of the detector, thereby evading AIGC detection.
R2BA solves 3 problems that existed in previous works:
Applicability to AIGC detectors. Based on the discovery that detectors vary in vulnerability to different post-processing techniques, we integrate 4 widely-used post-processing techniques to launch a comprehensive attack, ensuring all detectors are successfully attacked by R2BA, resulting in significant detection performance degradation.
@@
  • We propose R2BA, a novel realistic-like robust black-box adversarial attack. As the first method to evade AIGC detectors on natural images, R2BA provides significant practical value. It demonstrates applicability to black-box detectors, with impressive anti-detection performance, excellent image quality, and remarkable robustness.
  • We innovatively propose a post-processing fusion optimization method, which adjusts post-processing intensity to help post-processed AIGC images evade detection. Guided by the detector’s vulnerability to different post-processing, we use a stochastic particle swarm optimization algorithm with inertia decay to address the complex trade-off between adversariality and invisibility.
  • We conduct extensive experiments on the CNNSpot, GenImage, and DRCT datasets to validate the effective anti-detection performance of R2BA in evading SOTA GAN-based and diffusion-based AIGC detectors, including the commercial AIGC detector from Alibaba. Compared to existing attacks, R2BA achieves average improvements of 15%–72%, 21%–47%, and 26%–68% in adversariality, robustness and invisibility, respectively.
The remainder of this paper is organized as follows: Section II reviews related work. Section III presents the specific implementation of the proposed method. The experiments and analyses are described in Section IV. Finally, we provide a brief conclusion in Section V.
@@ II-A AI-generated Image Forgery Detection
AI-generated image forgery detection is actually a Fake-Real binary classification problem. Here, Fake refers to AIGC, which means the image is generated by AI, specifically through techniques like GANs or diffusion models. While Real refers to UGC, meaning that the image is a natural image or user-generated. The challenge of AIGC detection lies in extracting subtle forgery signals from AI-generated images and making accurate distinctions. According to different generation technologies, AIGC detection techniques can be divided into GAN-based and diffusion-based methods.
GAN-based forgery detection. Image generation based on GAN architecture[21],[22],[23] consists of generative and discriminative models. The goal is to train the generative model to match the real data distribution, making it difficult for the discriminative model to distinguish between real and AI-generated images. To prevent misuse, several GAN-based forgery detection techniques[10],[11],[24],[25],[26] have been proposed. These methods detect AI-generated images by extracting specific features and using data-driven approaches to learn differences across spatial, frequency, and statistical domains between AI-generated and natural images.
@@ II-B Adversarial Attacks for Anti-detection
Black-box attack. Black-box adversarial attacks do not require access to the inner workings of the model, and efficient attacks can be achieved by only obtaining the model output. There are three main types of black-box attacks: Algorithms based on gradient estimation[7] and agent models[30] are limited due to their reliance on fitting the target. When applied in practice, query-based methods might be the best solution. However, when dealing with the unpredictable decision boundaries caused by multi-class natural images[31], locating the decision boundary for a single adversarial example requires considerable effort. Some methods based on random pixel updates[16],[32] are effective in evading detection, but tend to introduce significant perturbations, which causes these adversarial examples to be directly classified as forged images by the human eye, even without a detector. Another method involving attribute editing[15], although optimizing multiple adversarial examples simultaneously, still face the challenge of selecting the appropriate attribute across the diverse natural images and requires suitable attribute editing tools.
Based on the above analysis, existing adversarial attacks for anti-detection have limitations in image classes (they only work on facial datasets) and detectors (primarily GAN-based detection). Additionally, they are often difficult to strike a balance between the anti-detection effectiveness and image quality. Moreover, in practical applications, real-world platform image post-processing may erase the adversarial perturbations, leading to the failure of attacks. To address these issues in existing works, we conduct comprehensive experiments on the security of GAN-based and diffusion-based AIGC detectors using natural image datasets with multiple classes, and thoroughly analyze and compare various anti-detection performance metrics of our method with existing methods in the original, robust and real-world scenarios.
Figure 2: Average performance degradation. The values represent the attack success rate of different post-processing against AIGC detectors. Darker colors indicate greater vulnerability of detectors. Attacks used include Gaussian blur (GB), JPEG compression (JG), Gaussian noise (GN), light spot (LS), and sequential fusion attacks (ALL means GB-JG-GN-LS).
@@ III-A Motivation
Traditional adversarial examples[6],[7],[16],[32] generated by gradient-based or random pixel modification attacks, are both meaningless and ineffective because their poor invisibility makes them difficult to apply in real-world scenarios. They can be easily identified as fake by the human eye, without the need for a detector. In contrast, adversarial examples created using real-world post-processing techniques appear realistic and match real natural image patterns. Furthermore, if a detector classifies such post-processed AIGC images as real, it suggests that the detector will experience similar performance degradation in real-world applications.
@@
Inspired by previous works[9],[18],[20], we select Gaussian blur, JPEG compression, Gaussian noise and light spot as attacks, which have been proven to mislead detectors through frequency, spatial, and statistical domains respectively. The pre-experiment we conduct in Fig. 2 further confirms that these post-processing methods lead to varying degrees of performance degradation in detectors. We use Gaussian blur with a radius of 5 and deviation of 1.5, JPEG compression with an intensity of 80, Gaussian noise with standard deviation of 7/255, and light spot with an intensity of 1.8 and random radius. The datasets[4],[24] and detectors[4],[10],[11],[24],[25] used in Fig. 2 are described in Section IV.
@@
The results in Fig. 2 demonstrate that detectors vary in vulnerability to different post-processing. For example, on the CNNSpot[24] dataset, the CNNSpot[24] detector is highly vulnerable to Gaussian noise but more robust to JPEG compression (GN=0.12, JG=0.00), while the FatFormer[11] detector exhibits the opposite behavior (GN=0.01, JG=0.23). In addition, the pre-experiment exhibits a counterintuitive result: The effectiveness of post-processing fusion attacks does not always exceed that of individual post-processing on different detectors. On the DRCT[4] dataset, the attack performance of Gaussian blur + JPEG compression on the FatFormer[11] is better than that of Gaussian blur alone (GB+JG=0.87, GB=0.18), whereas the opposite (GB+JG=0.02, GB=0.12) is true for the UnivFD[4]. As a result, optimizing the post-processing fusion intensity is challenging, but necessary.
@@
Therefore, considering that the detector is agnostic in real-world scenarios, we adopt a black-box attack to generate adversarial examples. To attack different types of detectors, we fuse these four post-processing and define the attack order as step 1-4 in Fig. 1. To solve the challenge of optimizing the fusion intensity, we use the detector’s fake probability as guidance and apply stochastic particle swarm optimization (PSO)[33] with inertia decay to optimize the fusion intensity. Compared to other optimization algorithms, such as evolutionary algorithms and simulated annealing, stochastic PSO with inertia decay outperforms them in effectively conducting global searches in high-dimensional spaces[34],[35]. This method can quickly locate the global decision boundary and simultaneously optimize fusion intensity to balance invisibility and adversariality. Based on the above-mentioned, we propose a post-processing fusion optimization-based, realistic-like robust black-box adversarial attack method, R2BA, to study the security of AIGC detectors in real-world applications.
III-B Visualization
@@
Figure 3: The pipeline of R2BA. It contains four modules and a main optimization process (Initialization-Optimization iteration-Selection). R2BA uses $N$ randomly generated post-processing parameters as particle positions, optimizing their velocities and positions to balance the invisibility and adversariality. A successful attack occurs when a particle crosses the detector’s decision boundary, and R2BA outputs the adversarial example with the highest SSIM value among the successful ones. The definitions of the notations here are the same as those in Table I.
@@
Component description. In Fig. 3, we illustrate the pipeline of R2BA. It mainly consists of four modules: the post-processing module, fitness calculation, velocity/position update, and the swarm optimization module. The decision boundary here refers to the fake threshold (0.5) of the detector, which typically classifies an image with a fake probability greater than 0.5 as fake, and less than 0.5 as real.
@@
1) Post-processing Module: The red circle represents an AIGC image, and the blue circles represent its post-processed images (adversarial particles). The particle positions represent the post-processing fusion intensity. We can balance invisibility and adversariality by optimizing the particle positions.
@@
2) Fitness Calculation: The post-processed image is fed into the detector to obtain its fake probability, which serves as the particle’s fitness. A smaller fitness value indicates a more effective adversarial example. The individual best of each particle corresponds to the smallest fitness in the optimization process, i.e., the lowest fake probability achieved so far. The global best is the particle with the smallest fitness among all individual bests. The global best particle is most likely to cross the decision boundary, as it has the lowest fake probability.
@@
3) Velocity / Position Update: Velocity $V_{i}^{k+1}$ updates are influenced by the particle’s individual best $p_{best}$, global best $P_{best}$, and inertia $w^{k}$. While the position of $p_{i}^{k+1}$ updates depend on the current position of $p_{i}^{k}$ and velocity $V_{i}^{k+1}$. Obviously, influenced by the individual best and global best, the particles will move towards positions with lower fake probability, continuously approaching the decision boundary.
@@
Optimization Iteration. In the first iteration, to cover as much of the example domain space as possible, R2BA randomly assigns $N$ post-processing fusion intensity values as the initial positions of $N$ particles. Stochastic generation helps quickly locate the detector’s most vulnerable decision boundary, enabling efficient global search and addressing the challenges posed by diverse decision boundaries in multi-class natural images. In subsequent iterations, taking the $k$-th iteration as an example, guided by fitness, the particle velocity gradually points towards the decision boundary, with the swarm becoming more concentrated. In the final (or successful) iteration, some particles cross the boundary, causing the detector to misclassify them as real images. R2BA then outputs the adversarial example with the highest SSIM among all the particles with successful attacks.
@@
The description above outlines a typical iterative process. However, in practice, AIGC images may be located near the decision boundary, causing the attack to succeed in the first iteration. If the images are far from the decision boundary, the attack may fail to cross it within the range defined by the post-processing intensity settings. Experiments show that the former case is more common, while the latter is relatively rare.